For a body starting from rest and moving with uniform acceleration, the distance travelled during the first t seconds and next t seconds is in the ratio:
A) 1:1 B) 1:2 C) 1:3 D)1:4 E) 1:9

Grade:12th pass

1 Answers

Eshan
5 years ago
Distance travelled in timetis

s=\dfrac{1}{2}at^2

Total distance travelled in time2t is

s'=\dfrac{1}{2}a(2t)^2=4s

Hence the distance travelled intseconds after the firsttseconds iss'-s=3s

henc teh required ratio is 1:3